Use the following data to answer the questions below.

P Q
$13.75 125
$13.20 129
$12.10 137
$11.55 145
$11.41 149

1.Using calculus to determine dQdPdQdP, construct a column which calculates the point-price elasticity for each (P,Q) combination. What is the point price elasticity of demand when P=$13.75P=$13.75?

Select one:

a. -0.861

b. -1.109

c. 125

d. -1.189

2.What is the point price elasticity of demand when P=$11.83P=$11.83?

Select one:

a. -0.835

b. -1.276

c. -0.144

d. 0.025
Answer #1

The regression results of this data provides a demand equation Q = 257.32 - 9.70P. This implies dQ/dP = -9.7. Slope of the demand curve is dQ/dP and so we have slope = -9.7.

1) Now When P is 13.75, Q is 125 so elasticity = -9.7 x 13.75 / 125 = -1.1. Hence select B)

2) When P is 11.83, Q is 142.25 so elasticity = -9.7 x 11.83 / 142.24= -0.8. Hence select A)